OpenStudy (anonymous):

Need Help ! What is the value of w? 5(9 − 3) = w • 9 – 5 • 3

OpenStudy (skullpatrol):

any ideas?

OpenStudy (anonymous):

no, i never did this before

OpenStudy (anonymous):

can you help me @skullpatrol ?

OpenStudy (skullpatrol):

Do know the distributive property?

OpenStudy (anonymous):

yes

OpenStudy (skullpatrol):

$$\Huge a(b-c)=ab-ac$$ $$\Huge 5(9 − 3) = w \cdot 9 – 5 \cdot 3$$

OpenStudy (anonymous):

ok , im going to try to work this out real fast

OpenStudy (skullpatrol):

Just look at it :) what must "a" equal?

OpenStudy (anonymous):

5

OpenStudy (skullpatrol):

Correct :)

OpenStudy (anonymous):

wait so thats the answe ? lol

OpenStudy (skullpatrol):

The value of w is 5.

OpenStudy (anonymous):

wow , thanks , that make me laugh because it was that hard(easy)

OpenStudy (skullpatrol):

math is that way sometimes :)

OpenStudy (skullpatrol):

No need to work things out, just step back and look at it.

OpenStudy (anonymous):

ok, can u help me out one more time so i can understand because that was just like WOW ! ?

OpenStudy (skullpatrol):

sure

OpenStudy (anonymous):

Which expressions show how 6 • 45 can be rewritten using the distributive property? 6 • 40 + 6 • 5 6 • 40 − 6 • 5 6 • 50 – 6 • 5 6 • 50 + 5

OpenStudy (skullpatrol):

Any ideas?

OpenStudy (anonymous):

2nd one ?

OpenStudy (skullpatrol):

no, there are two, but the second one is not right

OpenStudy (anonymous):

6 • 40 + 6 • 5

OpenStudy (skullpatrol):

yes

OpenStudy (skullpatrol):

one more please :)

OpenStudy (anonymous):

oh ok . ok

OpenStudy (anonymous):

6 • 50 – 6 • 5

OpenStudy (skullpatrol):

The distributive property is extremely important to be familiar and comfortable with.

OpenStudy (anonymous):

yea , this is the grade where i start learning it btw

OpenStudy (skullpatrol):

Multiplication is based on it.

OpenStudy (skullpatrol):

can you tell me the second one please?

OpenStudy (anonymous):

yea i see , im going to go watch a youtube video on this

OpenStudy (skullpatrol):

6 • 50 – 6 • 5

OpenStudy (anonymous):

yes

OpenStudy (skullpatrol):

$$\Huge 6 \cdot 45 = 6 \cdot 50 – 6 \cdot 5$$ $$\Huge a(b-c)=ab-ac$$

OpenStudy (anonymous):

so i was correct about 6 • 50 – 6 • 5

OpenStudy (anonymous):

ok, i see now

OpenStudy (skullpatrol):

Note how "a" is distributed as a multiplier of each of the terms ab and ac.

OpenStudy (anonymous):

THANKS !

OpenStudy (skullpatrol):

Thanks for trying so hard to learn something new :D

OpenStudy (anonymous):

:) yea its important to me aha